Can you please help...
The monthly income I, in dollars, from a new product is given by I(t)= 30,000-22,000e-0,008t where t is the time, in months, since the product was first put on the market. What was the monthly income after the 24th month? Round to the nearest dollar.

Answer by checkley77(12844)

IF THE (,) BEFORE THE 008 IS SUPPOSED TO BE A (.) THEN:
I(t)=30,000-22,000e-0,008t
I(t)=30,000-22,000*2.718-.008*24
I(t)=30,000-59,796-.192
I(t)=$-29,796